find the exact value of cos 105 degrees WITHOUT USING TABLES OR A CALCULATOR. be sure to show your work.
Answer by venugopalramana(3286)

find the exact value of cos 105 degrees WITHOUT USING TABLES OR A CALCULATOR. be sure to show your work.
COS(105)=COS(60+45)=COS(60)COS(45)-SIN(60)SIN(45)=(1/2)(1/SQRT2)-(SQRT3/2)(1/SQRT2)
=(1-SQRT(3))/2SQRT(2)=(1-1.732)/(2*1.414)=-0.732/2.828=-0.2588